OpenStudy (snowcrystal):

Help please I'm on a huge time crunch!!! I will fan and medal the most helpful person!!!!!!!

OpenStudy (snowcrystal):

What is the value of b in the triangle shown above? A.) -4inches B.) 4inches C.) negative or positive 4inches D.) no solution 

OpenStudy (anonymous):

Is the "24in. sq." pertaining to the area? or to that third side?

OpenStudy (snowcrystal):

the 3rd side

OpenStudy (***[isuru]***):

hmmm... i dont think its the 3rd side.. if it is the third side then u can use pythagoras theorem \[(AC) ^{2} = (AB)^{2} + (BC)^{2}\]\[24^{2} = (3b)^{2} + b^{2}\]\[24^{2} = 10b^{2}\]and obviously the answer is not given.. but its not the "no solution " since there is an value for "b" when u solve the above expression.. so the (24) .in. sq should be the area

OpenStudy (anonymous):

then that would be c^2 = a^2 + b^2 24^2 = 3b^2 + b^2 576 = 4b^2 All divided by 4 144 = b^2 \[\sqrt{144} = \sqrt{b ^{2}}\] 12 = b

OpenStudy (anonymous):

My answer comes with the pythagorean theorem, assuming it IS the 3rd side. if not, follow his answer^^^^^^

OpenStudy (***[isuru]***):

so... \[Area \ of \ a \ \triangle =\frac{ base \times height }{ 2 }\]\[= \frac{ AB \times BC }{ 2 }\]\[= \frac{ 3b \times b }{ 2 }\] since the area is 24 sq.in \[24 = \frac{ 3b \times b }{ 2 }\]\[48 = 3b \times b\]\[48 = 3b^{2}\]\[b^{2} = \frac{ 48 }{ 3 } = 16\]\[b = \pm \sqrt{16}\]which means \[b = +4 \ or \ -4\] but a length of a side of a triangle is always positive which means b = 4

OpenStudy (***[isuru]***):

got it :) ? OH! and btw ECSarabia u have done some miscalculations in ur simplifications... i think u should check it twice :)

OpenStudy (snowcrystal):

Yes I got it thanks a bunch!!!

OpenStudy (anonymous):

oh thanks. i've been a little off my game today. I think i'll stop "helping" because i have some stuff on my mind. Clouding my Arithmatic reasoning right now.

OpenStudy (snowcrystal):

it's okay to take a break every once in a while